Automotive air bags inflate when a sample of sodium azide, NaN , is very rapidly decomposed.




What mass of sodium azide is required to produce 2.6 ft (73.6 L) of nitrogen gas with a density of 1.25 g/L?

1
Expert's answer
2021-11-22T17:46:02-0500

Mass of N2 produced = density ×volume

= 1.25 g/L ×73.6 L

= 92 g


number of moles of N2 produced = mass/molar mass

= 92/28

= 3.286 mol


moles of NaN3 required = (2/3)×moles of N2

= (2/3)×3.286 mol

=2.19 mol


mass of NaN3 required= number of mol ×molar mass of NaN3

= 2.19×65

= 142.4 g